GAS-PHASE ELECTRONIC SPECTRA OF CARBON CHAINS

D. A. KIRKWOOD, H. LINNARTZ, T. MOTYLEWSKI, M. PACHKOV, M. TULEJ AND \underline{J. P. MAIER}, Institute of Physical Chemistry, University of Basel, Klingelbergstr. 80, CH-4056 Basel, Switzerland.

Electronic transitions of carbon chains and their ions are being measured in the gas-phase. The neutral species, such as C_{n}H n=6,8,10, have been observed by cavity ringdown absorption,^{a} whearas the carbon anions by a two photon detachment approach. Comparison is made with astronomical data on diffuse interstellar bands. The case of C_{7}^{−} is striking;^{b} the origin band of the A^{2}Π_{u}←X^{2}Π_{g} transition, as well as the strong vibronic transitions, coincide with DIBs. This provides compelling evidence for the first identification of a carrier.

aH. Linnartz, T. Motylewski and J. P. Maier, J. Chem. Phys., \underline{109}, 3819, (1998).

bM. Tulej, D. A. Kirkwood, M. Pachkov and J. P. Maier, Astrophys. J., \underline{506}, L69, (1998).